OpenStudy (aaaa):

Write the slope-intercept equation of the line perpendicular to y = 5/2 x - 2, which passes through the point (0, 2). A. y = 5x + 2 B. y = -2/5x + 2 C. y = 2x + 5 D. y = 1/2x + 5

OpenStudy (ash2326):

given line y=5/2x-2 standard form of line y=mx+c where m=slope comparing standard form to the given line we get slope m=5/2 we have to find the line which is perpendicular to this one let its slope be m1 m1*m=-1 m1*5/2=-1 m1=-2/5 so its equation is y=-2/5x+c we have to find c? it passes through (0,2) put x=0 y=2 to find c 2=0+c so c=2 y=-2/5x+2 option B